# Cycle products and efficient vectors in reciprocal matrices

### v1

@@ 2.3 Main result
Now, we may give the fundamental theorem on the cycle structure of reciprocal matrices and their efficient vectors. If $A\in\mathcal{PC}_{n}$ is consistent, then $\mathcal{E}(A)$ is the singleton generated by any column of $A,$ otherwise, since each column is in $\mathcal{E}(A)$ and not all columns are proportional, $\mathcal{E}(A)$ is not a singleton. For simplicity, we assume that $A$ is inconsistent.
@@
$\mathcal{E}(A)={\displaystyle\bigcup\limits_{C\in\pi_{<1}(A)}}\mathcal{\varepsilon}_{A}(\mathcal{C}).$ (4)
@@
$\mathcal{E}(A)={\displaystyle\bigcup\limits_{C\in\pi_{1}(A)}}\mathcal{\varepsilon}_{A}(\mathcal{C}).$ (5)
@@
Note that $\pi_{<1}(A)$ has no more than $\frac{(n-1)!}{2}$ elements and has less if there are Hamiltonian cycles from $A$ with product $1$. When $n=3,$ then $\frac{(n-1)!}{2}=1$ which means that $\mathcal{E}(A)=\varepsilon_{A}(\mathcal{C})$ for some cycle $\mathcal{C}\in$ $\pi_{1}(A),$ which is another way to see that $\mathcal{E}(A)$ is convex for $A\in\mathcal{PC}_{3}.$
@@
It is known that in general $\mathcal{E}(A)$ is not closed under component-wise geometric mean [18]. However, we note that each subset $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is, because it is defined by inequalities (2).
@@
Suppose $A\in\mathcal{PC}_{n}.$ If there exists a cycle that lies in $G(A,w)$ for every $w\in\mathcal{E}(A),$ then $\mathcal{E}(A)$ is closed under component-wise geometric mean.
@@
By Theorem 8, $\mathcal{E}(A)=\mathcal{\varepsilon}_{A}(\mathcal{C}_{3}),$ and, so, $\mathcal{E}(A)$ is convex. In particular, the Perron vector and the singular vector of $A$ are efficient for $A.$ Using (2), we can see that
@@
which can be verified to be contained in $\mathcal{\varepsilon}_{A}(\mathcal{C}_{3}),$ as expected. We observe that $A$ is not a simple perturbed consistent matrix, as it is nonsingular. Thus, this example illustrates that the convexity of $\mathcal{E}(A)$ may occur for other matrices $A$ than the simple perturbed consistent matrices.
3 $\mathcal{E}(A)$ uniquely determines $A$
Next we turn our attention to an interesting application of the fundamental theorem, other than the description of all efficient vectors.
Of course, $A\in\mathcal{PC}_{n}$ determines its efficient set $\mathcal{E}(A).$ We show that the converse is also so: $A\in\mathcal{PC}_{n}$ is uniquely determined by $\mathcal{E}(A).$ Our strategy is to show that, for every pair of distinct indices $i,j\in\{1,2,\ldots,n\},$ there is a vector $w\in\mathcal{E}(A)$ such that both $i\rightarrow j$ and $j\rightarrow i$ are edges in $G(A,w)$. This, means $w_{i}\geq a_{ij}w_{j}$ and $w_{j}\geq a_{ji}w_{i}.$ Since $a_{ij}a_{ji}=1,$ we must have $w_{i}=a_{ij}w_{j}$, i.e. $a_{ij}=\frac{w_{i}}{w_{j}},$ so that $a_{ij}$ (and, thus, $a_{ji}$) are determined by $w\in\mathcal{E}(A).$
Lemma 13
If $A\in\mathcal{PC}_{n}$ and $i,j\in\{1,\ldots,n\}$ with $i\neq j$, then there is a $\mathcal{C}\in\pi_{1}(A)$ that uses either the edge $i\rightarrow j$ or the edge $j\rightarrow i.$
Proof. Pick any cycle using $i\rightarrow j.$ If its product is $\leq 1,$ we are done. If not, the reverse of this cycle has the reciprocal product that is $\leq 1$ and uses the edge $j\rightarrow i.$
Lemma 14
Suppose that $A\in\mathcal{PC}_{n}$, that $w\in\mathbb{R}_{+}^{n}$ and that $i,j\in\{1,\ldots,n\}$ with $i\neq j$. If $\mathcal{C}$ is a Hamiltonian cycle in $G(A,w)$ that contains the edge $i\rightarrow j,$ then there is a vector $w^{\prime}\in\mathbb{R}_{+}^{n}$ such that $\mathcal{C}$ lies in $G(A,w^{\prime})$ and such that $G(A,w^{\prime})$ includes the additional edge $j\rightarrow i.$
Proof. Since $i\rightarrow j$ is an edge of $G(A,w),$ $w_{i}\geq a_{ij}w_{j}.$ Now define $w^{\prime}$ to agree with $w$ in all components except the $j$-th, and define $w_{j}^{\prime}=\frac{w_{i}}{a_{ij}}.$ Since $w_{j}^{\prime}\geq w_{j},$ $\mathcal{C}$ also lies in $G(A,w^{\prime}).$ Since $w^{\prime}$ satisfies $w_{i}^{\prime}=a_{ij}w_{j}^{\prime},$ we have $w_{j}^{\prime}=a_{ji}w_{i}^{\prime},$ verifying that $j\rightarrow i$ is also an edge of $G(A,w^{\prime}).$
Now we may apply our strategy to verify the following.
@@ 3 $\mathcal{E}(A)$ uniquely determines $A$
Let $A,B\in\mathcal{PC}_{n}$. If $\mathcal{E}(A)=\mathcal{E}(B)$ then $A=B.$
Proof. If $\mathcal{E}(A)$ is a singleton, then $A$ is consistent and the columns of $A$ are proportional to any vector in $\mathcal{E}(A)$ and are uniquely determined since all entries on the main diagonal of $A$ are $1$ Suppose that $\mathcal{E}(A)$ is not a singleton ($A$ is inconsistent) and $i,j\in\{1,2,\ldots,n\}$ with $i\neq j.$ By Lemma 13 and Theorem 8, there is a Hamiltonian cycle $\mathcal{C}$ containing one of the edges $i\rightarrow j$ or $j\rightarrow i$ and a vector $w\in\mathcal{E}(A)$ such that $G(A,w)$ contains $\mathcal{C}$. Then, Lemma 14 and Theorem 1 guarantee a $w^{\prime}\in\mathcal{E}(A)$ with $w_{i}^{\prime}=a_{ij}w_{j}^{\prime}$, implying that $a_{ij}=\frac{w_{i}^{\prime}}{w_{j}^{\prime}}.$ Thus, the $i,j$ entry of $A$ is determined. Since the pair $i,j$ is arbitrary, $A$ is fully determined.
4 Order reversals
Here we give some consequences of Theorem 8 regarding the existence of no order reversals in an efficient vector. Of course, if $A$ is consistent, no efficient vector for $A$ exhibits an order reversal, as $\frac{w_{i}}{w_{j}}=a_{ij}$ for all $i,j.$
Theorem 16
Suppose that $A\in\mathcal{PC}_{n}$ is inconsistent and $\mathcal{C}\in\pi_{1}(A).$ Then, there is a $w\in\varepsilon_{A}(\mathcal{C)}$ that exhibits no order reversals with entries of $A$ along $\mathcal{C}$ if and only if either there is an entry of $A$ $>1$ along $\mathcal{C}$ or the cycle product is $1$ Otherwise, there is a $w\in\varepsilon_{A}(\mathcal{C)}$ with at most $1$ order reversal along $\mathcal{C}$.
Proof. Suppose that there is a $w\in\varepsilon_{A}(\mathcal{C)}$ that exhibits no order reversals with entries of $A$ along $\mathcal{C}$ and all entries along $\mathcal{C}$ are $\leq 1,$ with at least one inequality strict. Then $\frac{w_{i}}{w_{j}}$ would be $\leq 1$ for the efficient vector $w$ and all $i,j$ along the cycle, with at least one inequality strict, implying
@@ 4 Order reversals
Suppose there is an entry $>1$ or all entries are $1$ along $\mathcal{C}$. For $\mathcal{C}$ as in (1), suppose, without loss of generality, that $a_{\gamma_{n}\gamma_{1}}\geq 1$, with $a_{\gamma_{n}\gamma_{1}}=1$ if and only if all entries along the cycle are $1$ Now, use the inequalities (2) and take the first $n-1$ of them to be equalities, so as to define $w,$ taking $w_{1}=1$ (Lemma 6). The last inequality (which is implied by the first $n-1$ inequalities) then insures that there is no order reversal at $i_{n},i_{1}$. In fact, $a_{\gamma_{n}\gamma_{1}}>1$ implies $\frac{w_{\gamma_{n}}}{w_{\gamma_{1}}}>1$. Also, $a_{\gamma_{n}\gamma_{1}}=1$ implies, by hypothesis, all entries of $A$ along the cycle equal to $1,$ and thus, by the $n-1$ equalities, $\frac{w_{\gamma_{j+1}}}{w_{\gamma_{j}}}=1,$ $11\leq j\leq n-$ 1 . Then, $\frac{w_{\gamma_{n}}}{w_{\gamma_{1}}}=1.$ There is no order reversal elsewhere along the cycle because of the equalities. The same construction insures at most one order reversal without the assumption of an entry $>1$ or all entries $1$ along $\mathcal{C}.$
@@ 5 Efficient vectors for column perturbed consistent matrices
Proof. Suppose that $B$is obtained from the consistent matrix $ww^{(-T)}$ by modifying, say, row and column $i.$ Then, for $D^{-1}=\operatorname*{diag}(w),$ we have that $DBD^{-1}$ has all entries equal $1$ except those in row and column $i.$ Then $DBD^{-1}$ is permutationally similar to a reciprocal matrix with all entries $1$ except those in the first row and column.
@@
We have that $\gamma$ can be one of the following permutations of $\{2,3,4,5\}$:
@@
Let $i,j\in\{2,\ldots,n\},$ $i\neq j.$ If $A\in\mathcal{PC}_{n}$ is as in (6) with $a_{1i}a_{j1}<1,$ then $\mathcal{\varepsilon}_{ij}(A)\subseteq\mathcal{E}(A).$ In particular, $\mathcal{\varepsilon}_{ij}(A)$ is the subset of $\mathcal{E}(A)$ that comes from the cycles of $\pi_{<1}(A)$ of the form $1\rightarrow i\rightarrow\cdots\rightarrow j\rightarrow 1$
Proof. Denote by $\mathcal{C}_{ij}$ the set of all Hamiltonian cycles $\mathcal{C}$ of the form $1\rightarrow i\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n-1}\rightarrow j\rightarrow 1$ (there are $(n-3)!$). Suppose that $a_{1i}a_{j1}<1,$ so that the product for any cycle $\mathcal{C}$ in $\mathcal{C}_{ij}$ is $<1.$ We have
@@
Proof. Let $\mathcal{C}$ be a Hamiltonian cycle in $A.$ By Lemma 20, $\mathcal{\varepsilon}_{ij}(A),$ with $(i,j)\in\mathcal{N}$, is contained in $\mathcal{E}(A),$ proving the inclusion $\supseteq$ in (7). On the other hand, by Theorem 8, if $w\in\mathcal{E}(A),$ then $w\in\varepsilon_{A}(\mathcal{C})$ for some $\mathcal{C}\in\pi_{<1}(A).$ Then, for $\mathcal{C}:1\rightarrow i\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n-1}\rightarrow j\rightarrow 1$, we have $a_{1i}a_{j1}<1.$ Since the cycle product for $\mathcal{C}$ in $A$ is $a_{1i}a_{j1},$ by Lemma 20, $\varepsilon_{A}(\mathcal{C})\subseteq\mathcal{\varepsilon}_{ij}(A)$.
The set $\mathcal{N}$ has at most $\frac{(n-1)(n-2)}{2}$ elements and has exactly this number if $a_{12},\ldots,a_{1n}$ are pairwise distinct.
@@
with $\mathcal{N}$ as in Theorem 21. Since, by Lemma 19, $\mathcal{\varepsilon}_{ij}(A)$ is convex then $S\mathcal{\varepsilon}_{ij}(A)$ is convex [21].
We finally show that the description of the efficient vectors for a simple perturbed consistent matrix given in [10] is a simple consequence of the results developed in this paper. Suppose that $A\in\mathcal{PC}_{n},$ $n>2,$ is as in (6), with $a_{12}=\cdots=a_{1,n-1}=1$ and $a_{1n}>1$ (if $a_{1n}=1,$ $A$ is consistent), which can be assumed for the purpose of studying the efficient vectors for a simple perturbed consistent matrix, by Lemma 2. Then, the Hamiltonian cycles whose products from $A$ are $<1$ are of the form $\mathcal{C}:1\rightarrow\ell\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n-1}\rightarrow n\rightarrow 1$, $\ell=2,\ldots,n-1.$ We have

### v3

@@ 2.3 Main result
Now, we may give the fundamental theorem on the cycle structure of reciprocal matrices and their efficient vectors. If $A\in\mathcal{PC}_{n}$ is consistent, then $\mathcal{E}(A)$ is the singleton generated by any column of $A,$ otherwise, since each column is in $\mathcal{E}(A)$ and not all columns are proportional, $\mathcal{E}(A)$ is not a singleton. We assume that $A$ is inconsistent.
@@
$\mathcal{E}(A)={\displaystyle\bigcup\limits_{\mathcal{C}\in\pi_{<1}(A)}}\mathcal{\varepsilon}_{A}(\mathcal{C}).$ (4)
@@
$\mathcal{E}(A)={\displaystyle\bigcup\limits_{\mathcal{C}\in\pi_{1}(A)}}\mathcal{\varepsilon}_{A}(\mathcal{C}).$ (5)
@@
Note that $\pi_{<1}(A)$ has no more than $\frac{(n-1)!}{2}$ elements and has less if there are Hamiltonian cycles from $A$ with product $1$. If $n=3,$ then $\frac{(n-1)!}{2}=1$ which means that $\mathcal{E}(A)=\varepsilon_{A}(\mathcal{C})$ for some cycle $\mathcal{C}\in$ $\pi_{1}(A)$ (in fact, $\mathcal{C}\in\pi_{<1}(A)$ if $A$ is inconsistent), which is another way to see that $\mathcal{E}(A)$ is convex for $A\in\mathcal{PC}_{3}.$
@@
It is known that in general $\mathcal{E}(A)$ is not closed under entry-wise geometric mean [18]. However, we note that each subset $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is (in fact, it is closed under entry-wise weighted geometric means), because it is defined by inequalities (2).
@@
Suppose $A\in\mathcal{PC}_{n}.$ If there exists a Hamiltonian cycle that lies in $G(A,w)$ for every $w\in\mathcal{E}(A),$ then $\mathcal{E}(A)$ is closed under entry-wise weighted geometric means.
@@
By Theorem 8, $\mathcal{E}(A)=\mathcal{\varepsilon}_{A}(\mathcal{C}_{3}),$ and, so, $\mathcal{E}(A)$ is convex. In particular, the Perron vector and the singular vector of $A$ are efficient for $A.$ Also, the set $\mathcal{E}(A)$ is closed under entry-wise weighted geometric means. Note that all the columns of $A$ lie in $\mathcal{\varepsilon}_{A}(\mathcal{C}_{3})$, as they are efficient for $A.$ Using (2), we can see that
@@
which can be verified to be contained in $\mathcal{\varepsilon}_{A}(\mathcal{C}_{3}),$ as expected. We observe that $A$ is not a simple perturbed consistent matrix, as it is nonsingular. Thus, this example illustrates that the convexity of $\mathcal{E}(A)$ may occur for other matrices $A$ than the simple perturbed consistent matrices. The matrix $A$ is also not double perturbed.
3 Order reversals
Here we give some consequences of Theorem 8 regarding the existence of order reversals in an efficient vector. Of course, if $A$ is consistent, no efficient vector for $A$ exhibits an order reversal, as $\frac{w_{i}}{w_{j}}=a_{ij}$ for all $i,j.$
@@ 3 Order reversals
Suppose that $A\in\mathcal{PC}_{n}$ is inconsistent and $\mathcal{C}\in\pi_{1}(A).$ Then, there is a $w\in\varepsilon_{A}(\mathcal{C)}$ that exhibits no order reversal with entries of $A$ along $\mathcal{C}$ if and only if either there is an entry of $A$ $>1$ along $\mathcal{C}$ or the cycle product is $1$ Otherwise, there is a $w\in\varepsilon_{A}(\mathcal{C)}$ with exactly $1$ order reversal along $\mathcal{C}$.
Proof. Suppose that there is a $w\in\varepsilon_{A}(\mathcal{C)}$ that exhibits no order reversal with entries of $A$ along $\mathcal{C}$ and all entries along $\mathcal{C}$ are $\leq 1,$ with at least one inequality being strict. Then $\frac{w_{i}}{w_{j}}$ would be $\leq 1$ for the efficient vector $w$ and all $i,j$ along the cycle, with at least one inequality being strict, implying
@@
Suppose there is an entry $>1$ or all entries are $1$ along $\mathcal{C}$. For $\mathcal{C}$ as in (1), suppose, without loss of generality, that $a_{\gamma_{n}\gamma_{1}}\geq 1$, with $a_{\gamma_{n}\gamma_{1}}=1$ if and only if all entries along the cycle are $1$ Now, use the inequalities (2) and take the first $n-1$ of them to be equalities, so as to define $w,$ taking $w_{1}=1$ (Lemma 6). The last inequality (which is implied by the first $n-1$ inequalities) then ensures that there is no order reversal at $\gamma_{n},\gamma_{1}$. In fact, $a_{\gamma_{n}\gamma_{1}}>1$ implies $\frac{w_{\gamma_{n}}}{w_{\gamma_{1}}}>1$. Also, $a_{\gamma_{n}\gamma_{1}}=1$ implies, by hypothesis, all entries of $A$ along the cycle equal to $1,$ and thus, by the $n-1$ equalities, $\frac{w_{\gamma_{j+1}}}{w_{\gamma_{j}}}=1,$ $11\leq j\leq n-$ 1 . Then, $\frac{w_{\gamma_{n}}}{w_{\gamma_{1}}}=1.$ There is no order reversal elsewhere along the cycle because of the equalities. The same construction ensures at most one order reversal without the assumption of an entry $>1$ or all entries $1$ along $\mathcal{C}.$
Theorem 13 concerns the existence of an efficient vector for $A$ with at most $1$ order reversal along a cycle $\mathcal{C}\in\pi_{1}(A).$ The vector may exhibit order reversals at positions in the matrix not along the cycle.
@@ 4 Efficient vectors for column perturbed consistent matrices
Proof. Suppose that $B$is obtained from the consistent matrix $ww^{(-T)}$ by modifying, say, row and column $i.$ Then, for $D^{-1}=\operatorname*{diag}(w),$ we have that $DBD^{-1}$ has all entries equal to $1,$ except those in row and column $i.$ Then $DBD^{-1}$ is permutationally similar to a reciprocal matrix with all entries equal to $1,$ except those in the first row and column.
@@
We have that $\gamma$ can be one of the following permutations of $\{2,3,4,5\}$ (we present them in a convenient order):
@@
Let $i,j\in\{2,\ldots,n\},$ $i\neq j.$ If $A\in\mathcal{PC}_{n}$ is as in (6) with $a_{1i}a_{j1}<1,$ then $\mathcal{\varepsilon}_{ij}(A)\subseteq\mathcal{E}(A).$ In particular, $\mathcal{\varepsilon}_{ij}(A)$ is the subset of $\mathcal{E}(A)$ that comes from the cycles in $\pi_{<1}(A)$ of the form $1\rightarrow i\rightarrow\cdots\rightarrow j\rightarrow 1$
Proof. Denote by $\mathcal{C}_{ij}$ the set of all Hamiltonian cycles $\mathcal{C}$ of the form $1\rightarrow i\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n-1}\rightarrow j\rightarrow 1$ (there are $(n-3)!$). Suppose that $a_{1i}a_{j1}<1,$ so that the product in $A$ for any cycle $\mathcal{C}$ in $\mathcal{C}_{ij}$ is $<1.$ We have
@@
Proof. Let $\mathcal{C}$ be a Hamiltonian cycle in $A.$ By Lemma 17, $\mathcal{\varepsilon}_{ij}(A),$ with $(i,j)\in\mathcal{N}$, is contained in $\mathcal{E}(A),$ proving the inclusion $\supseteq$ in (7). On the other hand, by Theorem 8, if $w\in\mathcal{E}(A),$ then $w\in\varepsilon_{A}(\mathcal{C})$ for some $\mathcal{C}\in\pi_{<1}(A).$ Then, for $\mathcal{C}:1\rightarrow i\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n-1}\rightarrow j\rightarrow 1$, the cycle product for $\mathcal{C}$ in $A$ is $a_{1i}a_{j1}<1.$ By Lemma 17, $\varepsilon_{A}(\mathcal{C})\subseteq\mathcal{\varepsilon}_{ij}(A)$.
The set $\mathcal{N}$ has at most $\frac{(n-1)(n-2)}{2}$ elements and has exactly this number if $a_{12},\ldots,a_{1n}$ are pairwise distinct. In fact, the number of elements in $\mathcal{N}$ is the number of pairs $(i,j)$ with $i,j\in\{2,\ldots,n\},$ $j>i,$ and $a_{1i}\neq a_{1j}.$
@@
with $\mathcal{N}$ as in the theorem. Since, by Lemma 16, $\mathcal{\varepsilon}_{ij}(A)$ is convex then $S\mathcal{\varepsilon}_{ij}(A)$ is convex [22].
We finally show that the description of the efficient vectors for a simple perturbed consistent matrix given in [10] is an easy consequence of the results developed in this paper. Suppose that $A\in\mathcal{PC}_{n},$ $n>2,$ is as in (6), with $a_{12}=\cdots=a_{1,n-1}=1$ and $a_{1n}>1$ (if $a_{1n}=1,$ $A$ is consistent), which can be assumed for the purpose of studying the efficient vectors for a simple perturbed consistent matrix, by Lemma 2. Then, the Hamiltonian cycles whose products from $A$ are $<1$ are of the form $\mathcal{C}:1\rightarrow\ell\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n-1}\rightarrow n\rightarrow 1$, $\ell=2,\ldots,n-1.$ We have
@@
Though the set of efficient vectors for a simple perturbed consistent matrix is convex, when the reciprocal matrix is double perturbed, that is, is obtained from a consistent matrix by changing two pairs of reciprocal entries, non-convexity may occur.
Example 20
Let
$A=\left[\begin{array}[c]{cccc}1&\frac{1}{3}&\frac{1}{2}&1\\ 3&1&1&1\\ 2&1&1&1\\ 1&1&1&1\end{array}\right].$
We have
$\mathcal{E}(A)=\mathcal{\varepsilon}_{23}(A)\cup\mathcal{\varepsilon}_{24}(A)\cup\mathcal{\varepsilon}_{34}(A),$
with
$\mathcal{\varepsilon}_{23}(A)$ =$=\left\{w\in\mathbb{R}_{+}^{4}:3w_{1}\geq w_{2}\geq w_{4}\geq w_{3}\geq 2w_{1}\right\},$
$\mathcal{\varepsilon}_{24}(A)$ =$=\left\{w\in\mathbb{R}_{+}^{4}:3w_{1}\geq w_{2}\geq w_{3}\geq w_{4}\geq w_{1}\right\},$
$\mathcal{\varepsilon}_{34}(A)$ =$=\left\{w\in\mathbb{R}_{+}^{4}:2w_{1}\geq w_{3}\geq w_{2}\geq w_{4}\geq w_{1}\right\}.$
We have
$u$ $=\left[\begin{array}[c]{cccc}1&3&3&3\end{array}\right]^{T}\in\mathcal{\varepsilon}_{23}(A),$
$v$ $=\left[\begin{array}[c]{cccc}1&1&2&1\end{array}\right]^{T}\in\mathcal{\varepsilon}_{34}(A).$
However,
$u+v=\left[\begin{array}[c]{cccc}2&4&5&4\end{array}\right]^{T}\notin\mathcal{E}(A).$
5 Conclusions
Examples of reciprocal matrices for which the set of efficient vectors is not convex are known, for example those for which the right Perron vector is not efficient. Here we have described the set of efficient vectors for a reciprocal matrix $A$ as a union of at most $\frac{(n-1)!}{2}$ convex sets. Each of these sets corresponds to a Hamiltonian cycle in $A$ whose product of the entries is less than $1$ (when the matrix is inconsistent). Our characterization allowed us to describe the set of efficient vectors for a reciprocal matrix obtained from a consistent matrix by modifying one column (row) as a union of at most $\frac{(n-1)(n-2)}{2}$ convex sets. In particular, the known characterization of the efficient vectors for a reciprocal matrix obtained from a consistent one by modifying one pair of reciprocal entries followed in an easy way. In this case the set of efficient vectors is convex. We also identified efficient vectors for $A$ with at most one order reversal at positions along the associated cycle.
The results obtained here may be helpful in a better understanding of the convexity of the set of efficient vectors for a reciprocal matrix, as well as of the existence of rank reversals in these vectors.
Acknowledgement 21
We thank the referees for the helpful comments.